# Structure of Classifier Boundaries: Case Study for a Naïve Bayes Classifier

### v2

@@ Abstract
Whether based on statistical models, training data or a combination, classifiers assign complex input data points to one of a small number of output categories. For a naïve Bayes classifier whose input space is a graph, we study the structure of the boundary, which comprises those points for which at least one neighbor is classified differently. The scientific setting is metagenomic assembly of DNA reads produced by Next Generation Sequencers (NGSs). The boundary is both large and complicated in structure. We introduce a new measure of uncertainty, Neighbor Similarity, that compares the result for an input point to the distribution of results for its neighbors. This measure not only tracks two inherent uncertainty measures for the Bayes classifier, but also can be implemented for classifiers without inherent measures of uncertainty.
Key words: classifier, Bayes classifier, uncertainty, DNA reads, metagenomics
@@ 1 Introduction
Classifiers are ubiquitous in today’s world. They are used in facial recognition, for military purposes, by autonomous vehicles, and, in this paper, as a step in assembling short DNA reads from multiple source genomes into longer “contigs.” Inputs include numerical data, character strings, images and sound clips. The underlying statistical methodologies range from maximum likelihood for mixture models to Bayes methods (as in this paper) to neural networks and deep learning based on labeled “training data.”
Typically, the input space for a classifier is both high-dimensional and enormous, while the output space is much smaller. Often, the input space is a graph, which is true in our case study. Then, insight into the classifier can be gained from understanding its boundary—those points in the input space with one or more neighbors that are classified differently. From the input data perspective, these points can be thought of as fragile, [^1] Our usage differs from that in epidemiology, where fragility is a property of clinical trials datasets that measures how many outcomes must change in order to alter statistical significance of the result. See, e.g., Baera et al., (2021). because changing them only ever so slightly changes the output. In our experiment, neighbors differ by one nucleotide pair (a single nucleotide polymorphism (SNP)), which can result from a sequencer error, another data quality problem, or natural polymorphism. And as we show below, quantifying how much a boundary point differs from its neighbors yields a measure of uncertainty that is computable for any classifier whose input space is a graph.
Understanding the boundary illuminates how the classifier works. Is the boundary large or small? Is its geometry simple or complex? How can it be located and explored? For our case study, the boundary in no way resembles boundaries for smooth functions, i.e., level sets. [^2] For instance, the level sets of a smooth function on a “continuous” space are smooth, lower-dimensional manifolds. By contrast, it is large, containing as much as 30% of the input space, not of lower dimension, and complex geometrically.
The remainder of this paper is organized in the following manner. In Section 2, we describe our experiment: the scientific setting, the data and the Bayes classifier, as well as introduce two surrogate uncertainty measures. Section 4 contains results for the DNA reads, together with techniques for exploring the boundary. Sections 5 and 6 contain discussion and conclusions, respectively.
@@ 2 Experimental Setting
Here we introduce the components of our case study. The illustrative problem we address is classifying short DNA sequences as having arisen from one of three candidate genomes, which is an essential step in software systems for metagenomic assembly.
@@ 2.1 Classifiers
In this paper, a classifier is a function $C$ from a finite input space $\mathcal{I}$ to a finite output space $\mathcal{O}$. Elements of $\mathcal{I}$ are termed inputs or data, and $C(x)\in\mathcal{O}$ is the decision or result for input $x$. Although it is not a logical necessity, there is little point in constructing or using a classifier unless $|\mathcal{I}|>>|\mathcal{O}|$, where $|S|$ is the cardinality of the set $S$. In our case study, 101 nucleotides with 5 possible values each yield $|\mathcal{I}|=5^{101}$, and three candidate genomes yield $|\mathcal{O}|=3$.
Although also not a necessity, we restrict attention to deterministic classifiers: for an input $x$, there is a unique, reproducible output $C(x)$. In our case study, $C$ is based on maximizing Bayesian posterior probabilities, and there are inherent measures of uncertainty to which we compare the surrogate measures proposed in Section 3.2. Mixture models that represent the effects of (discrete) latent variables are of the same ilk (Titterington et al.,, 1985). By contrast, deep learning and neural network models use a set $\mathcal{T}=\{t_{1},\dots,t_{m}\}\subseteq\mathcal{I}$ of labeled training data to determine tunable parameters of $C$ in such a way that $C(t_{j})$ is equal to a known value $D_{j}$ for each $j$.
We further assume that the input space $\mathcal{I}$ is an loopless, undirected graph, whose edgeset $E$ defines neighboring inputs: $x,y\in\mathcal{I}$ are neighbors if and only $x\neq y$ and $\{x,y\}\in E$. We denote by $N(x)=\{y:\{x,y\}\in E\}$ the set of neighbors, or neighborhood, of $x$. For ease of interpretation and because it is true in our case study, we further assume that $\mathcal{I}$ is connected: for any $x,y\in\mathcal{I}$, there is a path $(x_{1}=x,\dots,x_{k}=y)$ that connects them: $\{x_{j},x_{j+1}\}\in E$ for each $j$.
Our output space $\mathcal{O}$ consists of three virus genomes that give rise to simulated DNA reads produced by a NGS, and has no particular structure.
@@ 2.2 Mathematical Preliminaries
To establish notation, a DNA sequence $G$ is a character string chosen from the nucleotide (base) alphabet $\mathcal{N}=\{A,C,G,T\}$. At one extreme, $G$ may be an entire genome–for instance, a virus or a chromosome. At another, it may be a read generated by a NGS. Given a sequence $G$, its length is $|G|$; the $i^{\mathrm{th}}$ base in $G$ is $G(i)$; and the bases from location $i$ to location $j>i$ are $G(i:j)$.
We focus on triplets, whose distributions are 64-dimensional summaries of sequences. Other cases appear in the literature, especially quartets, also referred to as tetranucleotides (Pride et al.,, 2003; Teeling et al., 2004a, ; Teeling et al., 2004b, ). The triplet distribution $P_{3}(\cdot|G)$ of a sequence $G$ is defined as
$P_{3}(b_{1}b_{2}b_{3}|G)=\mathrm{Prob}\big{\{}G(k:k+2)=b_{1}b_{2}b_{3}\big{\}}$ (1)
for each choice of $b_{1}b_{2}b_{3}$, where $b_{1}$, $b_{2}$ and $b_{3}$ are elements of $\mathcal{N}$, and $k$ is chosen at random from $21,\dots,|G|-$ 2. An equivalent perspective is that of a second-order Markov chain (Karr et al.,, 2023). The information contained in $P_{3}(\cdot|G)$ is the same as that contained in the pair distribution $P_{2}(\cdot|G)$ and the $16\times 4$ transition matrix
$T_{3}((b_{1},b_{2}),b_{3}|G)=\mathrm{Prob}\big{\{}G(k+2)=b_{3}|G(k)=b_{1},G(k+1)=b_{2}\big{\}},$ (2)
@@ 2.3 The Three Genomes and the Reads Dataset
Our scientific context is reference-guided metagenomic assembly—piecing together fragments (reads) of DNA from multiple sources into longer sequences called contigs, as performed by MetaCompass (Cepeda et al.,, 2017). We focus on one conceptual step in assembly, namely, assigning reads to sources from a reference database. In our case study, there are three reference genomes: an adenovirus genome of length 34,125, downloaded with the read simulator Art, which we call Adeno; a SARS-CoV-2 genome of length 29,926 contained in a coronavirus dataset downloaded from National Center for Biotechnology Information (NCBI) in November, 2020, which we call COVID; and a SARS-CoV genome of length 29,751 from the same database, which we call SARS. Thus, $\mathcal{O}=\{\mbox{Adeno},\mbox{COVID},\mbox{SARS}\}$.
Table 9, in Appendix A, contains the triplet distributions for these three virus genomes. Measured by Hellinger distance, [^3] The definition is given in (7) below. these distributions are very different. The distances are 0.234 for Adeno/COVID, 0.125 for Adeno/SARS and 0.161 for COVID/SARS. To contextualize these values, the empirical 0.001 $p$-values for triplet distributions of genomes of the same size simulated from each distribution are 0.01941755 for adenovirus, 0.02094808 for COVID, and 0.02065539 for SARS. Therefore, classifiers based on triplet distributions are promising.
The read dataset is the same as in Karr et al., (2023). We employed the Mason_simulator software (Holtgrewe,, 2010) to simulate Illumina [^4] Illumina manufactures NGSs that employ an optical technology; see https://www.illumina.com/. NGS reads of length 101 from each genome, with approximate 6X coverage. The numbers of reads are 1966, 1996 and 1907, respectively; the total number of reads is 5869. The Mason_simulator introduces errors in the form of transpositions (SNPs), insertions, deletions and undetermined bases. [^5] These are cases in which the sequencer detects that a nucleotide is present, but is unable to determine whether it is A, C, G or T. Some DNA read datasets contain other “partial” classifications, for instance, a base that is one of two possibilities. The latter, following convention, appear in the simulated reads as “N” and must be accommodated in computations. Parameters of the Mason_simulator were set at default values.
@@ 2.4 The Bayes Classifier
The three likelihood functions, denoted by $L(\cdot|\mbox{Adeno})$, $L(\cdot|\mbox{COVID})$, and $L(\cdot|\mbox{SARS})$, calculated from the triplet distributions. To illustrate for adenovirus,
@@
×$\times T_{3}((R(2),R(3)),R(4)|\mbox{Adeno})\times\dots$
×$\times T_{3}\bigg{(}\big{(}R(|R|-2),R(|R|-1)\big{)},R(|R|)|\mbox{Adeno}\bigg{)},$
where $P_{2}(\cdot|A)$ is the pair distribution defined by analogy with (1), and $T_{3}$ is given by (2). In (3), we have ignored Ns for simplicity; when they are present, they lead to sums over all possible bases.
@@ 3.1 The Boundary
=$\mathcal{B}=\{R\in\mathcal{I}:C(R^{\prime})\neq C(R)\mbox{ for some }R^{\prime}\in N(R)\}.$ (5)
@@
Note that whenever two reads are classified differently, any path in $\mathcal{I}$ connecting them must contain at least two (neighboring) boundary points. Below we focus on Hamming paths, which are the shortest. Given reads $R\neq R^{\prime}$ with Hamming distance $k$—they differ in $k$ of 101 locations, a Hamming path from $R$ to $R^{\prime}$ simply replaces one differing nucleotide in $R$ at a time by the corresponding nucleotide in $R^{\prime}$. Therefore, each successive pair on the path are neighbors. Such a path $\big{(}R_{0}=R,\dots,R_{k}=R^{\prime}\big{)}$ has length $k+1$; there are $k!$ of them. The “simplest” path simply moves left to right in the character strings. Because $C(R)\neq C(R^{\prime})$, there must be at least one value of $j$ for which $C(R_{j})\neq C(R_{j+1})$, in which case, the neighbors $R_{j}$ and $R_{j+1}$ both belong to $\mathcal{B}$.
To provide initial numerical evidence, Table 1 contains the confusion matrix for the Bayes classifier $C$, which is computable because we know the source of each read, and shows that $C$ performs well, albeit not spectacularly. Looking only at the columns, Table 1 identifies $3,867,933=1933\times 2001$ pairs, the first of which is classified as Adeno and the other as COVID; $3,740,355=1933\times 1935$ pairs, the first of which is classified as Adeno and the other as SARS; $3,871,935=2001\times 1935$ pairs, the first of which is classified as COVID and the other as SARS. Any conclusion about the seemingly large size of these numbers should be tempered by remembering that $|\mathcal{I}|=5^{101}$.
@@ 3.2 Surrogate Uncertainty Measures
The Bayes classifier $C$ possesses two inherent and related measures of uncertainty, the maximum posterior likelihood and the entropy of the posterior distribution. [^6] The entropy of a probability distribution $P$ on a finite set $\mathcal{S}$ is $H(P)=-\sum_{s\in\mathcal{S}}p(s)\log p(s)$, with the convention that $\times-\infty=0$ 0. Entropy is minimized by distributions concentrated at a single point and maximized at the uniform distribution on $\mathcal{S}$, with maximizing value $\log(|\mathcal{S}|)$. Other classifiers may lack such measures, although there might be informal measures of “confidence.” For instance, partition models often characterize confidence in terms of the homogeneity of terminal nodes, which is a useful but not fully principled measure, since it does not result from a statistical model.
We now introduce three surrogate measures that are computable for any classifier operating on the same input space $\mathcal{I}$. Two of them, for a read $R$, are functions of the distribution on $\mathcal{O}$ of $\{C(R^{\prime}):R^{\prime}\in N(R)\}$.
@@
The second measure, Neighbor Similarity, quantifies the extent to which the neighbors of a read have the same decision it does. In symbols, given a read $R$,
@@
where $H$ denotes Hellinger distance (Nikulin,, 2001), $q_{R}$ is the degenerate probability distribution on $\mathcal{O}$ concentrated on $C(R)$, and $Q_{N(R)}$ is the probability distribution on $\mathcal{O}$ of $\{C(R^{\prime}):R^{\prime}\in N(R)\}$. The Hellinger distance between distributions $p$ and $q$ on a finite set $A$ is given by
Hellinger distance is a metric, and takes values between zero (if and only if $p=q$) and one (when $p$ and $q$ have disjoint supports). In symbols,
=$\mathcal{B}=\{R\in\mathcal{I}:\mbox{NS}(R)<1\}.$
Unlike BS, NS takes account only of how many neighbors differ, not their values. The almost inescapable interpretation of $\mbox{NS}(R)$ is uncertainty regarding the classification of $R$. If $\mbox{NS}(R)=1$, all neighbors of $R$ have the same decision as $R$, and, intuitively, we would be more certain of the decision for $R$. At the other extreme, if $\mbox{NS}(R)$ were zero, [^7] Which does not occur in our case study. changing any single nucleotide of $R$ changes the decision, in which case we would be highly uncertain about $C(R)$.
Both BS and NS can be calculated for any classifier whose input space has the graph structure described in Section 2, and their values returned along with the classifier decision. The only but not trivial burden is computational: the classifier must be run on all neighbors of the input. See Section 5 for further discussion.
A third surrogate measure is equally intuitive but more speculative, because we lack a demonstrably feasible method for calculating it. It is the Distance from the Boundary: $\mbox{DB}(R)$ is the minimum value of $k$ for which there exists a path $(R_{0}=R,\dots,R_{k})$ in $\mathcal{I}$ such that $R_{k}\in\mathcal{B}$. Then, the interpretation is that the larger $\mbox{DB}(R)$, the greater the certainty regarding $C(R)$, because the more $R$ would have to be “mutated” in order to change $C(R)$. It is easy to obtain upper bounds on $C(R)$: for any other read $R^{\prime}$ with $C(R^{\prime})\neq C(R)$, $\mbox{DB}(R)$ is at most the Hamming distance between $R$ and $R^{\prime}$. The unresolved problem is exact computation of $\mbox{DB}(R)$. Also, the ubiquity of the boundary suggests that DB may lack discriminatory power not already present in BS and NS.

### v3

# Structure of Classifier Boundaries: Case Study for a Bayes Classifier

@@ Abstract
Classifiers assign complex input data points to one of a small number of output categories. For a Bayes classifier whose input space is a graph, we study the structure of the boundary, which comprises those points for which at least one neighbor is classified differently. The scientific setting is assignment of DNA reads produced by Next Generation Sequencers (NGSs) to candidate source genomes. The boundary is both large and complicated in structure. We introduce a new measure of uncertainty, Neighbor Similarity, that compares the result for an input point to the distribution of results for its neighbors. This measure not only tracks two inherent uncertainty measures for the Bayes classifier, but also can be implemented for classifiers without inherent measures of uncertainty.
@@ 1 Introduction
Classifiers are ubiquitous in today’s world. They are used in facial recognition, for military purposes, by autonomous vehicles, and, in this paper, to classify short DNA reads from multiple potential source genomes. Classifier inputs include numerical data, character strings, images and sound clips. The underlying statistical methodologies range from maximum likelihood for mixture models to Bayes methods (as in this paper) to neural networks and deep learning based on labeled “training data.”
Typically, the input space for a classifier is both high-dimensional and enormous, while the output space is much smaller. When the input space is a graph, which is true in our case study, insight into the classifier can be gained from understanding its boundary—those points in the input space with one or more neighbors that are classified differently. From the input data perspective, these boundary points can be thought of as fragile, because changing them only ever so slightly changes the output. (This usage of the word “fragile” differs from that in epidemiology, where fragility is a property of clinical trials datasets that measures how many outcomes must change in order to alter statistical significance of the result. See, e.g., Baera et al., (2021).) In our experiment, neighbors differ by one nucleotide (a single nucleotide polymorphism (SNP)), which in real sequence data can result from a sequencer error, another data quality problem, or natural polymorphism. Quantifying how much a boundary point differs from its neighbors yields a measure of uncertainty that is computable for any classifier whose input space is a graph.
Understanding the boundary illuminates how the classifier works. Is the boundary large or small? Is its geometry simple or complex? How can it be located and explored? For our case study, the boundary is large, containing as much as 30% of the input space, not of lower dimension, and complex geometrically. By contrast, boundaries for a “smooth” function on a “continuous” space—namely, the level sets—are continuous, lower-dimensional manifolds.
The remainder of this paper is organized in the following manner. In Section 2, we describe our experiment: the scientific setting, the data and the Bayes classifier, as well as introduce two surrogate uncertainty measures. Section 4 contains results for the DNA reads, together with techniques for exploring the boundary. Sections 5 and 6 contain discussion and conclusions.
@@ 2 Experimental Setting
Here we introduce the components of our case study. The problem we address is classifying short DNA sequences generated by so-called next generation sequencers as having arisen from one of three candidate genomes. The dataset we employ is synthetic, so that ground truth is known. The broader scientific context is reference-guided metagenomic assembly—piecing together fragments (reads) of DNA from multiple potential sources into longer sequences called contigs, as performed, for example, by MetaCompass (Cepeda et al.,, 2017), with the assistance of a (reference) database of sequenced candidate genomes. The role of the reference database is to reduce the computational burden of creating the contigs, which is formulated as finding an Eulerian path in a de Bruijn graph (Compeau et al.,, 2023). The specific problem setting that led to this study involved samples from human digestive and reproductive tracts (Porter and Karr,, 2024).
@@ 2.1 Classifiers
In this paper, a classifier is a function $C$ from a finite input space $\mathcal{I}$ to a finite output space $\mathcal{O}$. Elements of $\mathcal{I}$ are termed inputs or data, and $C(x)\in\mathcal{O}$ is the decision or result for input $x$. Although it is not a logical necessity, there is little point in constructing or using a classifier unless $|\mathcal{I}|>>|\mathcal{O}|$, where $|\mathcal{S}|$ is the cardinality of the set $\mathcal{S}$. In our case study, 101 nucleotides in each DNA read with five possible values each (see below for an explanation of the fifth value) yield $|\mathcal{I}|=5^{101}$, and three candidate genomes yield $|\mathcal{O}|=3$.
We restrict attention to deterministic classifiers: for an input $x$, there is a unique, reproducible output $C(x)$. In our case study, $C$ is based on maximizing Bayesian posterior probabilities, and there are inherent measures of uncertainty to which we can compare the surrogate measures proposed in Section 3.2. Mixture models that represent the effects of (discrete) latent variables are of the same ilk (Titterington et al.,, 1985). By contrast, deep learning and neural network models use a set $\mathcal{T}=\{t_{1},\dots,t_{m}\}\subseteq\mathcal{I}$ of labeled training data to determine tunable parameters of $C$ in such a way that $C(t_{j})$ is equal to a known value $D_{j}$ for each $j$.
We further assume that the input space $\mathcal{I}$ is an loopless, undirected graph, whose edgeset $E$ defines neighboring inputs: $x,y\in\mathcal{I}$ are neighbors if and only $x\neq y$ and $\{x,y\}\in E$. We denote by $\mathcal{N}(x)=\{y:\{x,y\}\in E\}$ the set of neighbors, or neighborhood, of $x$. For ease of interpretation and because it is true in our case study, we further assume that $\mathcal{I}$ is connected: for any $x,y\in\mathcal{I}$, there is a path $(x_{1}=x,\dots,x_{k}=y)$ that connects them: $\{x_{j},x_{j+1}\}\in E$ for each $j$.
Our output space $\mathcal{O}$ consists of three virus genomes that give rise to (simulated) DNA reads produced by a particular NGS.
@@ 2.2 Mathematical Preliminaries
To establish notation, a DNA sequence $S$ is a character string chosen from the nucleotide (base) alphabet $\{A,C,G,T\}$, where A is adenine, C is cytosine, G is guanine and T is thiamine. At one extreme, $S$ may be an entire genome–for instance, a virus or a chromosome. At another, it may be a read generated by a NGS. Given a sequence $S$, its length is $|S|$; the $i^{\mathrm{th}}$ base in $S$ is $S(i)$; and the bases from location $i$ to location $j>i$ are $S(i:j)$.
We focus on triplets, whose distributions are 64-dimensional summaries of sequences. That triplets suffice in our case study is discussed below. Additional, more general, discussion appears in Karr et al., (2023). And, of course, triplets are fundamental biologically because they encode amino acids. Other cases appear in the literature, especially quartets, also referred to as tetranucleotides (Pride et al.,, 2003; Teeling et al.,, 2004). The triplet distribution $P_{3}(\cdot|S)$ of a sequence $S$ is defined as
$P_{3}(b_{1}b_{2}b_{3}|S)=\mathrm{Prob}\big\{S(k:k+2)=b_{1}b_{2}b_{3}\big\}$ (1)
for each choice of $b_{1}b_{2}b_{3}$, where $b_{1}$, $b_{2}$ and $b_{3}$ are elements of $\{A,C,G,T\}$, and $k$ is chosen at random from $1,\dots,|S|-2$. An equivalent perspective is that of a second-order Markov chain (Karr et al.,, 2023). The information contained in $P_{3}(\cdot|S)$ is the same as that contained in the pair distribution $P_{2}(\cdot|S)$ and the $16\times 4$ transition matrix
$T_{3}((b_{1},b_{2}),b_{3}|S)=\mathrm{Prob}\big\{S(k+2)=b_{3}|S(k)=b_{1},S(k+1)=b_{2}\big\},$ (2)
@@ 2.3 The Three Genomes and the Reads Dataset
In our case study, there are three reference genomes: an adenovirus genome of length 34,125, downloaded with the read simulator Art, which we call Adeno; a SARS-CoV-2 genome of length 29,926 contained in a coronavirus dataset downloaded from National Center for Biotechnology Information (NCBI) in November, 2020, which we call COVID; and a SARS-CoV genome of length 29,751 from the same database, which we call SARS. These are typical lengths for virus genomes. Thus, $\mathcal{O}=\{\mbox{Adeno},\mbox{COVID},\mbox{SARS}\}$.
Table 11, in Appendix A, contains the triplet distributions for these three virus genomes. Measured by Hellinger distance, whose definition is given in (6) below, these distributions are very different. The distances are 0.234 for Adeno/COVID, 0.125 for Adeno/SARS and 0.161 for COVID/SARS. To contextualize these values, the .999-quantiles for Hellinger distances of triplet distributions in samples of the same size simulated from each genome are 0.01941755 for adenovirus, 0.02094808 for COVID, and 0.02065539 for SARS. Therefore, none of the triplet distributions is remotely likely to have arisen from another genome, so that classifiers based on triplet distributions.
The read dataset is the same as in Karr et al., (2023). We employed the Mason_simulator software (Holtgrewe,, 2010) to simulate Illumina NGS reads of length 101 from each genome, with approximate 6X coverage. (Illumina manufactures NGSs that employ an optical technology; see https://www.illumina.com/.) The numbers of reads are 1966, 1996 and 1907, respectively, which sum to 5869. The Mason_simulator introduces errors in the form of transpositions (SNPs), insertions, deletions and undetermined bases. The latter are cases in which the sequencer detects that a nucleotide is present, but is unable to determine whether it is A, C, G or T. Following convention, these appear in the simulated reads as “N,” and must be accommodated in computations. Some other DNA read datasets contain additional “partial” classifications, for instance, a base that is one of two possibilities.
Parameters of the Mason_simulator were set at default values.
@@ 2.4 The Bayes Classifier
The three likelihood functions, denoted by $L(\cdot|\mbox{Adeno})$, $L(\cdot|\mbox{COVID})$, and $L(\cdot|\mbox{SARS})$, are calculated from the triplet distributions. To illustrate for adenovirus,
@@
$\times T_{3}((R(2),R(3)),R(4)|\mbox{Adeno})\times\dots$
$\times T_{3}\bigg(\big(R(|R|-2),R(|R|-1)\big),R(|R|)|\mbox{Adeno}\bigg),$
where $P_{2}(\cdot|A)$ is the pair distribution defined by analogy with (1), and $T_{3}$ is given by (2). In (3), we have ignored Ns for simplicity; when they are present, they lead to sums over all possible bases that could have been the N.
@@ 3.1 The Boundary
$\mathcal{B}=\{R\in\mathcal{I}:C(R^{\prime})\neq C(R)\mbox{ for some }R^{\prime}\in\mathcal{N}(R)\}.$ (5)
@@
Whenever two reads are classified differently, any path in $\mathcal{I}$ connecting them must contain at least two (neighboring) boundary points. Below we focus on Hamming paths, which are the shortest. Given reads $R\neq R^{\prime}$ with Hamming distance $k$ (They differ in $k$ of 101 locations.), a Hamming path from $R$ to $R^{\prime}$ simply replaces one differing nucleotide in $R$ at a time by the corresponding nucleotide in $R^{\prime}$. Therefore, each successive pair on the path are neighbors. Such a path $\big(R_{0}=R,\dots,R_{k}=R^{\prime}\big)$ has length $k+1$; there are $k!$ of them. The “simplest” of these paths simply moves left to right in the character strings. Because $C(R)\neq C(R^{\prime})$, there must be at least one value of $j$ for which $C(R_{j})\neq C(R_{j+1})$, in which case, the neighbors $R_{j}$ and $R_{j+1}$ both belong to $\mathcal{B}$.
To provide initial numerical evidence, Table 1 contains the confusion matrix for the Bayes classifier $C$, which is computable because we know the source of each read. The classifier performs well, albeit not spectacularly. (See Section 4.6.2 for brief discussion of other classifiers.) Looking only at the columns, Table 1 identifies $3,867,933=1933\times 2001$ pairs, the first of which is classified as Adeno and the other as COVID; $3,740,355=1933\times 1935$ pairs, the first of which is classified as Adeno and the other as SARS; $3,871,935=2001\times 1935$ pairs, the first of which is classified as COVID and the other as SARS. Any conclusion about the seemingly large size of these numbers should be tempered by remembering that $|\mathcal{I}|=5^{101}$.
@@ 3.2 Surrogate Uncertainty Measures
The Bayes classifier $C$ possesses two inherent and related measures of uncertainty, the maximum posterior likelihood and the entropy of the posterior distribution. (The entropy of a probability distribution $P$ on a finite set $\mathcal{S}$ is $H(P)=-\sum_{s\in\mathcal{S}}p(s)\log p(s)$, with the convention that $0\times-\infty=0$. Entropy is minimized by distributions concentrated at a single point and maximized at the uniform distribution on $\mathcal{S}$, with maximizing value $\log(|\mathcal{S}|)$.) Other classifiers may lack such measures, although there might be informal measures of “confidence.” For instance, partition models often characterize confidence in terms of the homogeneity of terminal nodes, which is a useful but not fully principled measure, since it does not result from a statistical model.
We now introduce three surrogate uncertainty measures that are computable for any classifier operating on the same input space $\mathcal{I}$. Two of them, for a read $R$, are functions of the distribution on $\mathcal{O}$ of $\{C(R^{\prime}):R^{\prime}\in\mathcal{N}(R)\}$.
@@
The second measure, Neighbor Similarity, quantifies the extent to which the neighbors of a read have the same decision it does, using Hellinger distance (Nikulin,, 2001). The Hellinger distance between distributions $p$ and $q$ on a finite set $A$ is given by
Hellinger distance is a metric, and takes values between zero (if and only if $p=q$) and one (when $p$ and $q$ have disjoint supports).
Neighbor Similarity is then defined for a read $R$ as
@@
where $H$ denotes Hellinger distance, $q_{R}$ is the degenerate probability distribution on $\mathcal{O}$ concentrated on $C(R)$, and $q_{N(R)}$ is the probability distribution on $\mathcal{O}$ of $\{C(R^{\prime}):R^{\prime}\in\mathcal{N}(R)\}$. Therefore,
$\mathcal{B}=\{R\in\mathcal{I}:\mbox{NS}(R)<1\}.$
Unlike BS, NS takes account only of how many neighbors differ, not their values. Our principal proposed interpretation of $\mbox{NS}(R)$ is as uncertainty regarding the classification of $R$. If $\mbox{NS}(R)=1$, all neighbors of $R$ have the same decision as $R$, and, intuitively, we would be more certain of the decision for $R$. At the other extreme, if $\mbox{NS}(R)$ were zero, changing any single nucleotide of $R$ changes the decision, in which case we would be highly uncertain about $C(R)$. This extreme case does not occur in our case study.
However, as discussed further in Section 4.6.1, we must distinguish certainty from correctness. In some regions of $\mathcal{I}$, the Neighbor Similarity may be high because “everything” is classified incorrectly.
Both BS and NS can be calculated for any classifier whose input space has the graph structure described in Section 2, and their values returned along with the classifier decision. The burden is computational: the classifier must be run on all neighbors of the input. See Section 5 for further discussion.
A third surrogate measure is more speculative, because we lack a demonstrably efficient method for calculating it. It is the Distance from the Boundary: $\mbox{DB}(R)$ is the minimum value of $k$ for which there exists a path $(R_{0}=R,\dots,R_{k})$ in $\mathcal{I}$ such that $R_{k}\in\mathcal{B}$. Then, the interpretation is that the larger $\mbox{DB}(R)$, the greater the certainty regarding $C(R)$, because the more $R$ would have to be “mutated” in order to change $C(R)$. It is easy to obtain upper bounds on $C(R)$: for any other read $R^{\prime}$ with $C(R^{\prime})\neq C(R)$, $\mbox{DB}(R)$ is at most the Hamming distance between $R$ and $R^{\prime}$. The unresolved problem is exact computation of $\mbox{DB}(R)$. Also, the ubiquity of the boundary suggests that DB may lack discriminatory power not already present in BS and NS.